OpenStudy (anonymous):

solve for u 6+(6/u-1)=(5/u+3)

OpenStudy (anonymous):

6+6/u-1=5/u+3 multiply both sides by u 6u+6-u=5+3u 5u-6=5+3u 2u=11 u=11/2

OpenStudy (anonymous):

or 5.5 if you want a decimal

OpenStudy (anonymous):

wait the donominater are polynomials so it 6 /(u-1) and 5/(u+3)

OpenStudy (anonymous):

6+6/u-1=5/u+3 multiply both sides by u 6u+6-u=5+3u 5u-6=5+3u 2u=11 u=11/2

OpenStudy (anonymous):

dont forget medals

OpenStudy (anonymous):

ok, thats more annoying 6+6/(u-1)=5/(u+3) multiply by both denominators 6*(u+3)*(u-1)+6*(u+3)=5*(u-1)

OpenStudy (anonymous):

expanding and multiplying, 6*(u^2+2u-3)+6u+18=5u-5 6u^2+12u-18+6u+18=5u-5 6u^2+18u=5u-5

OpenStudy (anonymous):

get it all to one side 6u^2+18u=5u-5 6u^2+13u+5=0

OpenStudy (anonymous):

then use the quadratic equation to get your roots

OpenStudy (anonymous):

i forgot that how does it go

OpenStudy (anonymous):

\[(-b \pm \sqrt{b^2-4AC})/2A\] where the equation is of the form Ax^2+Bx+C=0

OpenStudy (amistre64):

\[6+\frac{6}{u-1}=\frac{5}{u+3}\text{ ; /6}\] \[1+\frac{1}{u-1}=\frac{5}{6(u+3)}\text{ ; /5}\] \[\frac{1}{5}+\frac{1}{5(u-1)}=\frac{1}{6(u+3)}\text{ ; common denom it}\] \[\frac{u}{5(u-1)}=\frac{1}{6(u+3)}\text{ ; just a thought lol}\]

OpenStudy (anonymous):

amis, those dont make the equation any easier

OpenStudy (amistre64):

i dunno; looks pretty easy to me :) \[\frac{u}{5(u-1)}=\frac{1}{6(u+3)}\text{ ; /u}\] \[\frac{1}{5u(u-1)}=\frac{1}{6u(u+3)}\text{ ; /u}\] \[\frac{1}{5u^2-5u}=\frac{1}{6u^2+18u}\text{ }\] 6u^2 +18u = 5u^2 -5u u^2 +23u = 0 u(u+23) = 0 u = 0 or -23 ; did i miss it?

OpenStudy (amistre64):

i missed it; 5u is bad bad bad

OpenStudy (amistre64):

6u^2 +18u = 5u -5 6u^2 +13u + 5

OpenStudy (anonymous):

you should get it down to my equation and have to use the quadratic equation anyway, your way just takes longer

OpenStudy (amistre64):

have fun with it; i gots ta run ;)